hi sir,how can i minimze the time for solving yough problens.is there any trick which can save my time.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

hi palash

cracking IIT JEE problem is always not use

there are many tricks to solve tough problem

but a single trick u can't use for every problem

tricks changes accroding to question that u can learn by solving question

more question u will solve u will learn more tricks u will develope ur own..

for example

if (a+x+x2)6n= a0 +a1x+a2x2+........

then find a0+a3+.......

here if u will solve it will take lot of time

but u can take n=1/3 then u can find a0+a3+....... (only for objectives question)